6. Joy’s Frozen Yogurt shops have enjoyed rapid growth in northeastern states in recent years. From
the analysis of Joy’s various outlets, it was found that the demand curve follows this pattern:
Q = 200 - 300P + 120I + 65T - 250Ac + 400Aj


a. What would be the effect of a $5,000 increase in the competitor’s advertising expenditure?
Illustrate the effect on the outlet’s demand curve.
b. What would Joy’s advertising expenditure have to be to counteract this effect?
1
Expert's answer
2018-10-26T10:40:09-0400

Q = 200 - 300P + 120I + 65T - 250Ac + 400Aj

a. The effect of a $5,000 increase in the competitor’s advertising expenditure will cause the decrease in Joy's demand of 250*5,000 = $1,250,000.

b. Joy’s advertising expenditure would have to increase by 1,250,000/400 = $3,125 to counteract this effect.